Sam can paint a house in 5 days by himself, while Betty can paint the same house in 20 days. How long would it take Sam and Betty to paint the house together?
Answer by richwmiller(17219)

1/a+1/b=1/x
1/5+1/20=1/x
20/100+5/100=1/x
25/100=1/x
x=100/25
x=4 days working together
check
x/a+x/b=1
x/5+x/20=1
4/5+4/20=1
0.8+0.2=1
1=1
ok
codewt